Home‎ > ‎Crew 133‎ > ‎


Cap Com from XO Gartrelle Crew #133

Daily report on Grenhab Ops 1/6/2014

Heater set to 40 degrees at 9AM
Heater set to 75 degrees at 5PM

Fan Operating Normally

No watering done today. Watering is planned for tomorrow AM. Will check newly planted seedlings for moisture and if dry will add a small amount of water to the bottom of the tray of cubes per documentation.

Per your communication tonight the seedlings have been moved to the upstairs part of the hab where it is warmest.. However I am concerned they will not receive enough direct sunlight here during the day. Please provide your comments or suggestions on this. 

Greenhab photos planned for tomorrow afternoon at 3PM  

Thermostat Readings from 5PM today:

                                    Current                              Min                                Max
Temperature              70 Degrees F                  37 Degrees F                   84 Degrees F
Humidity                        17%                                 13%                                 17%      

In response to your question about needing a description of the planting activity, the following was submitted yesterday during Comms. Please provide any specific questions you have about what we have done. 
At 3:45 Sweet Peppers, Hot peppers, Mint Lettuce and Ithaka Lettuce seeds were planted in Root Riot Cubes. This work was done in the main hab. Twelve cubes with approximately two seeds each were planted for each of the four varieties. (The lettuce seeds were extremely hard to handle). The cubes were covered and returned to the green hab by 4:40. Crewman Howell ably assisted in the work. Excess seeds were returned to the can in the main hab.